A = bh, subject to the constraint b - h = 7.
Use the constraint to express b in terms of h.
A in terms of h.
please answer with the steps of how to do it..
thank you in advance

Answer by rothauserc(4718)

b in terms of h
:
b - h = 7
:
b = h + 7
:
A in terms of h
:
A = bh
:
substitue for b
:
A = (h + 7)h
:
A = h^2 + 7h
: